# Probabilistic Pontryagin’s Maximum Principle for Continuous-Time Model-Based Reinforcement Learning

### v2

@@ IV-B Two-point boundary value problem
For each realization $\boldsymbol{\theta}_{i}$, the initial state $\boldsymbol{x}(t_{0};\boldsymbol{\theta}_{i})=\boldsymbol{x}_{0}$ and the final co-state $\boldsymbol{\lambda}(t_{f};\boldsymbol{\theta}_{i})=\nabla_{\boldsymbol{x}}\Phi\bigl{(}\boldsymbol{x}(t_{f};\boldsymbol{\theta}_{i})\bigr{)}$ are shared. This gives rise to a two-point boundary value problem (TPBVP) that satisfies the coupled differential equations $\dot{\boldsymbol{x}}(t;\boldsymbol{\theta}_{i})=\boldsymbol{f}\bigl{(}\boldsymbol{x}(t;\boldsymbol{\theta}_{i}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}_{i})\bigr{)}$ and $\dot{\boldsymbol{\lambda}}(t;\boldsymbol{\theta}_{i})=-\nabla_{\boldsymbol{x}}\mathcal{H}(\boldsymbol{x}(t;\boldsymbol{\theta}_{i}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}_{i}),\boldsymbol{u}(t),t;\boldsymbol{\theta}_{i})$, where the initial value of the co-state $\boldsymbol{\lambda}(t_{0};\boldsymbol{\theta}_{i})$ is unknown. Due to the continuous-time formulation, numerical ODE solvers with adaptive step-size solvers can be leveraged to improve stability and control the numerical error.
We solve this using the forward method [17], which guesses an initial co-state $\boldsymbol{\lambda}(t_{0};\boldsymbol{\theta}_{i})$, integrates the states and co-states forward, and adjusts the guess based on the termination errors $\boldsymbol{\lambda}(t_{f};\boldsymbol{\theta}_{i})-\nabla_{\boldsymbol{x}}\Phi\bigl{(}\boldsymbol{x}(t_{f};\boldsymbol{\theta}_{i})\bigr{)}$. Root-finding methods (e.g., Newton’s method) or least-squares optimization (e.g., Levenberg-Marquardt) refine the guess iteratively. For robustness with non-linear dynamics for long time horizons, we employ the multiple shooting formulation [17]. This method breaks down the long time horizon into multiple shorter intervals and enforces continuity conditions between these, to improve convergence and stability.
@@ IV-C Model predictive control
The TPBVP describes an open-loop trajectory optimization problem. To obtain a closed-loop control, principles from non-linear MPC can be leveraged by solving the trajectory optimization at every time point under a moving horizon of $t\in[t_{0},t_{0}+H]$ [17]. The optimized control function is then applied until a new system measurement becomes available, and the optimization process is repeated.
@@ V Experimental Results
We demonstrate our approach on three classical control problems. Code implementing the experiments is available at https://github.com/DavidLeeftink/probabilistic-pontryagin-control.
@@ V-A Van der Pol stabilization
$\dot{x}_{1}$ $=\mu\left(x_{2}-\frac{1}{3}x_{1}^{3}-x_{1}\right)$
$\dot{x}_{2}$ $=-x_{1}+u$
@@
We formulate an offline RL task by generating a a dataset of $n=25$ trajectories (length $T=10$) with the Van der Pol system, whose dynamics are assumed unknown, under random periodic control inputs (Schroeder sweeps [19]) from uniformly sampled initial conditions. We then use this data to model the dynamics probabilistically, by training an ensemble of neural ODEs which capture the epistemic uncertainty. The resulting probabilistic dynamical model is then used to construct a controller by either following the probabilistic Pontryagin approaches or by directly optimizing the mean cost.
To explicitly investigate the effect of epistemic uncertainty modeling, we compare our approach using both deterministic and probabilistic dynamical models trained on identical datasets. We benchmark the mean Hamiltonian (PMP-Mean $\mathcal{H}$) and mean posterior (PMP-Mean $u^{*}$) formulations against direct optimization methods: Sequential Quadratic Programming (SQP), Broyden–Fletcher–Goldfarb–Shanno (BFGS) [20], improved Cross-Entropy Method (iCEM) [21, 22], and Adam [23]. We provide further experimental details in the Appendix, Table III.
TABLE I: Comparison of optimization methods for Van der Pol stabilization with true and learned dynamics (mean cost $\pm$ std. dev. over 15 datasets).
@@
Table I summarizes the performance in terms of accumulated cost over 15 repetitions. The indirect PMP approach achieves improved cost performance over its direct counterpart, under both the deterministic and probabilistic dynamical model. The lowest cost is achieved under the probabilistic Pontryagin approaches, where the mean optimal control approximation slightly outperforms the mean Hamiltonian approach, suggesting that explicitly modeling the epistemic uncertainty improves performance.
@@ V-B Cart pole swing up
TABLE II: Comparison of optimization methods for the Cart pole control problem with true and learned dynamics (mean cost $\pm$ std. dev. over 15 datasets).
@@
$\ddot{x}$ $=\frac{lm\sin(\theta)\dot{\theta}^{2}+u+m\,g\cos(\theta)\sin(\theta)}{M+m(1-\cos^{2}(\theta))}$
[…]
We design an offline RL task in similar fashion to the Van der Pol experiment, where the dynamics of the cart pole are assumed to be unknown and have to be learned from observed trajectories. We generate $n=50$ trajectories using random controls and uniformly sampled initial conditions, which serve as data. This results in sparse observations near the goal state. We compare the performance of the previously mentioned trajectory optimization methods on deterministic and probabilistic neural ODEs. Further experimental details are described in the Appendix, Table III.
Figure 1: Comparison of optimization methods for the cart pole control task with deterministic and probabilistic dynamical models. The boxplot shows the distribution of cost over 15 datasets, including the median, interquartile range, and outliers. Due to its high cost average, the Adam optimizer is not visualized.
Table II and Figure 1 present the trial costs averaged over 15 datasets, comparing deterministic and probabilistic dynamical models across various optimization methods. Under limited available data, adopting the probabilistic dynamical model improves performance for most optimizers, emphasizing the role of capturing epistemic uncertainty in learning-based control tasks. The mean Hamiltonian formulation achieves the lowest mean and median trial costs out of the considered optimization methods, outperforming direct methods. In contrast, the mean posterior approximation, theoretically suited only for convex Hamiltonians, performs worse with higher variability and average cost.
@@ V-C Duffing oscillator stabilization
$\dot{x}_{1}$ $=x_{2}$
$\dot{x}_{2}$ $=-\delta x_{2}-\alpha x_{1}-\beta x_{1}^{3}+\gamma u$
@@
In the online RL setting, one starts from a single random-policy trial, and iteratively trains the probabilistic dynamical model, updates the controller, and augments the dataset after each trial. We compare the mean Hamiltonian and mean posterior methods to the state-of-the-art PETS algorithm [12], which combines probabilistic dynamical models with MPC and the (i)CEM optimizer. We adapt PETS to the continuous-time setting using probabilistic neural ODEs as dynamical models. Further experimental details are described in the Appendix, Table III.
Figure 2: Comparison of mean Hamiltonian PMP approach with PETS, for online RL of a Duffing oscillator (cost $\pm$ std. dev. over 10 independent repetitions).
@@ VI Conclusion and discussion
We have proposed a probabilistic extension of PMP that incorporates epistemic uncertainty by minimizing the mean of Hamiltonian functions with respect to a posterior distribution over system dynamics. This approach enables the application of classical PMP techniques to probabilistic dynamical models, which commonly arise in learning-based control and RL settings, particularly under limited data.
Our numerical implementation combines forward multiple shooting with moving horizon optimization. This method indirectly optimizes the state and co-state trajectories, implicitly defining continuous-time optimal control functions. This naturally aligns with the continuous-time nature of many real-world control systems, while still only requiring discrete-time observations. Although we adopted neural ODE ensembles as a probabilistic dynamical model, the proposed framework generalizes easily to other probabilistic dynamical models, including Bayesian neural ODEs or systems with known functional forms but uncertain parameters.
Our experiments demonstrate the practical effectiveness of our approach in offline RL tasks under limited data availability, including the Van der Pol oscillator and a cart pole system. Comparisons against direct trajectory optimization methods, such as SQP and (i)CEM, show that our proposed mean Hamiltonian formulation achieves improved trial costs. Our results further underline the advantage of explicitly modeling the epistemic uncertainty: Under identical neural architectures, the probabilistic ensemble model consistently outperforms its deterministic counterpart, indicating that capturing epistemic uncertainty can lead to improved control performance. Additionally, in online RL settings, such as the stabilization of a Duffing oscillator, the proposed approach achieves competitive trial cost and data efficiency, comparable to the state-of-the-art MBRL method PETS.
@@
Future theoretical work could explore extending the optimality conditions to bounded control sets, a common setting for many practical control problems including the experiments considered in this work. From a practical perspective, our probabilistic PMP approach is especially suited for applications requiring precise control under uncertainty on slower timescales, such as personalized treatment planning in healthcare or process control, where state constraints like dosage or safety limits often arise. The PMP framework can be extended by incorporating the state constraints in the Hamiltonian via additional co-state variables.
@@ ACKNOWLEDGMENTS
We would like to thank Dr. Yuzhen Qin for his valuable feedback and insightful comments during the review of this manuscript.
@@
TABLE III: System parameters and control cost matrices used in the experiments.
System Parameters $Q$ $Q_{f}$ $R$ $\sigma$ $\mathcal{U}$ $x_{0}$ $x^{*}$ $t_{0}$ $t_{f}$ $\Delta t$ (MPC) $H$ Solver
Van der Pol $\mu=1.5$ $\boldsymbol{I}$ $\boldsymbol{I}$ $0.5$ $0.01$ $[-2,2]$ $(1,1)^{\top}$ $(0,0)^{\top}$ 00 $10$ $0.05$ $3$ dopri8
Cart pole $l$ $=m=M=1$ $g$ $=9.81$ $(1,1,0.1,0.1)^{\top}$ $(1,5,1,1)^{\top}$ $50.0$ $[-20,20]$ $(0,0,0,0)^{\top}$ $(1,\pi,0,0)^{\top}$ 00 $5$ $0.02$ $1$ dopri8
Duffing $\alpha$ $=-1,\quad\beta=2$ $\delta$ $=0.2,\quad\gamma=1$ $5\cdot\boldsymbol{I}$ $5\cdot\boldsymbol{I}$ $1$ $0.01$ $[-2,2]$ $(1.5,1)^{\top}$ $(0,0)^{\top}$ 00 $5$ $0.05$ $2$ dopri8
@@ VI-A Experimental details
For all experiments, cubic interpolation is used to construct a continuous-time control function from a discretized control vector, and a 5th-order Dormand-Prince solver is used to solve the states (and co-states for PMP) [27]. The following settings were used for the respective experiments:
@@
Following the approach described in [12], we consider an ensemble size of $M=5$, with each neural ODE consisting of two hidden layers with 32 units, resulting in parameters $\boldsymbol{\theta}\in\mathbb{R}^{1248}$. Our indirect methods use $S=4$ shooting segments. All trajectory optimization methods employ moving horizon optimization (horizon $H=3$) and 15 iterations per timestep.
@@
The dynamical model architecture is consistent with the previous Van der Pol setup, using ensemble neural ODEs ($\boldsymbol{\theta}\in\mathbb{R}^{1376}$) with identical control interpolation and ODE solver. Here, we use a shorter MPC horizon of $H=1$ with 25 iterations per time step, and two shooting segments ($S=2$) to ensure computational efficiency.
@@
We adopt an ensemble of neural ODEs with $M=5$ as a probabilistic dynamical model, consistent with the previous van der Pol setup ($\boldsymbol{\theta}\in\mathbb{R}^{1248}$) with identical control interpolation and ODE solver. We employ MPC with horizon $H=2$ and adjusted integration steps suitable for the Duffing oscillator’s faster time scale, and use $S=5$ shooting segments for the indirect methods.
The system parameters and control cost matrices used in all experiments are described in Table III.

### v3

@@ IV-B Two-point boundary value problem
For each realization $\boldsymbol{\theta}_{i}$, the initial state $\boldsymbol{x}_{\boldsymbol{\theta}_{i}}(t_{0})=\boldsymbol{x}_{0}$ and the final co-state $\boldsymbol{\lambda}_{\boldsymbol{\theta}_{i}}(t_{f})=\nabla_{\boldsymbol{x}}\Phi\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}_{i}}(t_{f})\bigr{)}$ are shared. This gives rise to a two-point boundary value problem (TPBVP) that satisfies the coupled differential equations $\dot{\boldsymbol{x}}_{\boldsymbol{\theta}_{i}}(t)=\boldsymbol{f}_{\boldsymbol{\theta}_{i}}\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}_{i}}(t),\boldsymbol{\lambda}_{\boldsymbol{\theta}_{i}}(t)\bigr{)}$ and $\dot{\boldsymbol{\lambda}}_{\boldsymbol{\theta}_{i}}(t)=-\nabla_{\boldsymbol{x}}\mathcal{H}_{\boldsymbol{\theta}_{i}}(\boldsymbol{x}_{\boldsymbol{\theta}_{i}}(t),\boldsymbol{\lambda}_{\boldsymbol{\theta}_{i}}(t),\boldsymbol{u}(t),t)$, where the initial value of the co-state $\boldsymbol{\lambda}_{\boldsymbol{\theta}_{i}}(t_{0})$ is unknown. Due to the continuous-time formulation, numerical ODE solvers with adaptive step-size solvers can be leveraged to improve stability and control the numerical error.
We solve this using the forward shooting method [21], which guesses an initial co-state $\boldsymbol{\lambda}_{\boldsymbol{\theta}_{i}}(t_{0})$, integrates the states and co-states forward, and adjusts the guess based on the terminal errors $\boldsymbol{\lambda}_{\boldsymbol{\theta}_{i}}(t_{f})-\nabla_{\boldsymbol{x}}\Phi\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}_{i}}(t_{f})\bigr{)}$. Root-finding methods (e.g., Newton’s method) or least-squares optimization (e.g., Levenberg-Marquardt) refine the guess iteratively. For robustness with non-linear dynamics for long time horizons, we employ the multiple shooting formulation [21]. This method breaks down the long time horizon into multiple shorter intervals and enforces continuity conditions between these, to improve convergence and stability.
@@ IV-C Model predictive control
The TPBVP describes an open-loop trajectory optimization problem. To obtain a closed-loop control, principles from non-linear MPC can be leveraged by solving the trajectory optimization at every time point under a moving horizon of $t\in[t_{0},t_{0}+H]$ [21]. The optimized control function is then applied until a new system measurement becomes available, and the optimization process is repeated. By using a finite prediction horizon, the control resulting from MPC is not necessarily optimal because it does not account for how future control actions will be re-optimized based on new measurements. Thus, this approach trades off theoretical optimality for computational tractability and the ability to react to real-time disturbances.
@@ V Experimental Results
We demonstrate our approach on three classical control problems, with all system parameters, control cost matrices and simulation parameters detailed in the Appendix, Table III. For each task, we adopt a standard quadratic cost functional of the form
$C(\boldsymbol{u})=(\boldsymbol{x}(t_{f})-\boldsymbol{x}^{*})^{\top}Q_{f}(\boldsymbol{x}(t_{f})-\boldsymbol{x}^{*})$
$+\int_{t_{0}}^{t_{f}}\left((\boldsymbol{x}(t)-\boldsymbol{x}^{*})^{\top}Q(\boldsymbol{x}(t)-\boldsymbol{x}^{*})+\boldsymbol{u}(t)^{\top}R\boldsymbol{u}(t)\right)dt$
where $Q$ and $Q_{f}$ are positive semi-definite state-cost matrices, $R$ is a positive definite control-cost matrix, and $\boldsymbol{x}^{*}$ is the target state. The code implementing the experiments is available at https://github.com/DavidLeeftink/probabilistic-pontryagin-control.
@@ V-A Van der Pol stabilization
$\dot{x}_{1}$ $=\mu\left(x_{2}-\frac{1}{3}x_{1}^{3}-x_{1}\right),$
$\dot{x}_{2}$ $=-x_{1}+u,$
@@
We formulate an offline RL task where we generate a dataset by simulating the Van der Pol system for $n=25$ trajectories of length $T=10$. These simulations are driven by random periodic control inputs (Schroeder sweeps [23]) and initialized from uniformly sampled states. In this offline setting, the system’s underlying dynamics are assumed to be unknown. We then use this data to train a probabilistic dynamics model — an ensemble of neural ODEs — to capture the epistemic uncertainty. Finally, the learned model is used to construct a controller, allowing a direct comparison between the probabilistic Pontryagin approaches and direct optimization approaches of the mean cost.
To explicitly investigate the effect of epistemic uncertainty modeling, we compare our approach using both deterministic and probabilistic dynamics models trained on identical datasets. We benchmark the mean Hamiltonian (PMP-Mean $\mathcal{H}$) and mean posterior (PMP-Mean $u^{*}$) formulations against direct optimization methods: Sequential Quadratic Programming (SQP), Broyden–Fletcher–Goldfarb–Shanno (BFGS) [24], improved Cross-Entropy Method (iCEM) [25, 26], and Adam [27]. We provide further experimental details in the Appendix, Table III.
TABLE I: Comparison of optimization methods for Van der Pol stabilization with true and learned dynamics (mean cost $\pm$ std. dev. over 15 repetitions).
@@
Table I summarizes the performance in terms of accumulated cost over 15 repetitions. The indirect PMP approach achieves improved cost performance over its direct counterpart, under both the deterministic and probabilistic dynamics model. The lowest cost is achieved under the probabilistic Pontryagin approaches, where the mean optimal control approximation slightly outperforms the mean Hamiltonian approach, suggesting that explicitly modeling the epistemic uncertainty improves performance.
@@ V-B Cart pole swing up
TABLE II: Comparison of optimization methods for the Cart pole control problem with true and learned dynamics (mean cost $\pm$ std. dev. over 15 repetitions).
@@
$\ddot{x}$ $=\frac{lm\sin(\theta)\dot{\theta}^{2}+u+m\,g\cos(\theta)\sin(\theta)}{M+m(1-\cos^{2}(\theta))},$
$\ddot{\theta}$ $=-\frac{lm\cos(\theta)\sin(\theta)\dot{\theta}^{2}+u\cos\theta+(M+m)\,g\sin(\theta)}{lM+lm(1-\cos^{2}(\theta))}.$
Here, $x$ represents the horizontal position of the cart, $\theta$ is the angle of the pole (with $\theta=\pi$ corresponding to the upright, balanced position), and the control input $u$ is the horizontal force applied to the cart. The physical parameters are the pole’s half-length $l$, the pole’s mass $m$, the cart’s mass $M$, and the acceleration due to gravity $g$.
We design an offline RL task in similar fashion to the Van der Pol experiment, where the dynamics of the cart pole are assumed to be unknown and have to be learned from observed trajectories. We generate $n=50$ trajectories of length $T=5$ using random controls and uniformly sampled initial conditions, which serve as data. This results in sparse observations near the goal state. We compare the performance of the previously mentioned trajectory optimization methods on deterministic and probabilistic neural ODEs. Further experimental details are described in the Appendix, Table III.
Figure 3: Comparison of optimization methods for the cart pole control task with deterministic and probabilistic dynamics models. The boxplot shows the distribution of cost over 15 datasets, including the median, interquartile range, and outliers. Due to its high cost average, the Adam optimizer is not visualized.
Table II and Figure 3 present the trial costs averaged over 15 repetitions, comparing deterministic and probabilistic dynamics models across various optimization methods. Under limited available data, adopting the probabilistic dynamics model improves performance for most optimizers, emphasizing the role of capturing epistemic uncertainty in learning-based control tasks. The mean Hamiltonian formulation achieves the lowest mean and median trial costs out of the considered optimization methods, outperforming direct methods. The mean posterior approximation, theoretically suited only for convex Hamiltonians, performs worse than its counterpart with higher variability and average cost.
@@ V-C Duffing oscillator stabilization
$\dot{x}_{1}$ $=x_{2},$
$\dot{x}_{2}$ $=-\delta x_{2}-\alpha x_{1}-\beta x_{1}^{3}+\gamma u.$
@@
In the online RL setting, one starts from a single random-policy trial, and iteratively trains the probabilistic dynamics model, updates the controller, and augments the dataset after each trial. We compare the mean Hamiltonian and mean posterior methods to the state-of-the-art PETS algorithm [16], which combines probabilistic dynamics models with MPC and the (i)CEM optimizer. We adapt PETS to the continuous-time setting using probabilistic neural ODEs as dynamics models. Further experimental details are described in the Appendix, Table III.
Figure 4: Comparison of online model-based RL methods for the Duffing oscillator stabilization task (cost $\pm$ std. dev. over 10 independent repetitions).
@@ VI Conclusion and discussion
In this work, we have introduced a novel framework for MBRL that addresses the critical challenge of planning under epistemic uncertainty. Our approach translates a probabilistic interpretation of PMP, established in optimal control theory [10, 11], into a practical algorithm for systems with learned dynamics. The core of our method is the principle of minimizing the mean Hamiltonian, which we compute over an ensemble of models representing the posterior distribution over the system dynamics that arises from limited data.
We developed a scalable numerical method, combining forward multiple shooting with moving horizon optimization, to make this principled approach tractable for deep, probabilistic models. A key feature of our framework is that it optimizes trajectories in continuous time—naturally aligning with the physics of many real-world systems—while still only requiring standard discrete-time observations for model learning. While we adopted neural ODE ensembles as a probabilistic dynamics model, the proposed principle generalizes to other probabilistic dynamics models, including Bayesian neural ODEs or systems with known functional forms but uncertain parameters.
Our experiments demonstrate the practical effectiveness of our approach in offline RL tasks under limited data availability, including the Van der Pol oscillator and a cart pole system. Comparisons against direct trajectory optimization methods, such as SQP and (i)CEM, show that our proposed mean Hamiltonian formulation achieves reduced trial costs. Our results further underline the advantage of explicitly modeling the epistemic uncertainty: under identical neural architectures, the probabilistic ensemble model consistently outperforms its deterministic counterpart, indicating that capturing epistemic uncertainty can lead to improved control performance. Additionally, in online RL settings, such as the stabilization of a Duffing oscillator, the proposed approach achieves competitive trial cost and data efficiency, comparable to the state-of-the-art MBRL method PETS.
@@
Future theoretical work could explore extending the optimality conditions to problems with state constraints, which commonly arises in safety critical application. One possible avenue to realize this is by incorporating safety constraints in the Hamiltonian via additional co-state variables. From a practical perspective, our probabilistic PMP approach is especially suited for applications requiring precise control under uncertainty on slower timescales, such as personalized treatment planning in healthcare or process control.
@@ ACKNOWLEDGMENTS
We would like to thank Dr. Yuzhen Qin for his valuable feedback and insightful comments during the review of this manuscript. We are also grateful to the anonymous reviewers for their constructive suggestions, which helped to significantly improve the final version of the paper.
@@
TABLE III: System parameters, control cost matrices and simulation parameters used in the experiments.
System Parameters $Q$ $Q_{f}$ $R$ $\sigma$ $\mathcal{U}$ $_{0} xx^{*}$ $t_{0}$ $t_{f}$ $\Delta t$ (MPC) $H$
Van der Pol $\mu=1.5$ $\mathbb{I}$ $\mathbb{I}$ $0.5$ $0.01$ [2,2][-2,2] $(1,1)^{\top}$ $(0,0)^{\top}$ 0 $10$ $0.05$ 33
Cart pole $l$ $=m=M=1$ $g$ $=9.81$ $\mathbb{I}(1,1,0.1,0.1)^{\top}$ $\mathbb{I}(1,5,1,1)^{\top}$ $0.05$ 0 [20,20][-20,20] $(0,0,0,0)^{\top}$ $(1,\pi,0,0)^{\top}$ 0 55 $0.02$ 11
Duffing $\alpha$ $=-1,\quad\beta=2$ $\delta$ $=0.2,\quad\gamma=1$ $5\cdot\mathbb{I}$ $5\cdot\mathbb{I}$ 11 $0.01$ [2,2][-2,2] $(1.5,1)^{\top}$ $(0,0)^{\top}$ 0 55 $0.05$ 22
@@ VI-A Experimental details
For all experiments, cubic interpolation is used to construct a continuous-time control function from a discretized control vector, and a 5th-order Dormand-Prince solver is used to solve predicted states (and co-states for PMP) [31] for the learned dynamics model. An overview of all the system parameters, control cost matrices and simulation parameters used in the experiments is presented in Table III.
The following settings were used for the respective experiments:
@@
Following the approach described in [16], we consider an ensemble size of $M=5$ following common practice in model-based RL, with each neural ODE consisting of two hidden layers with 32 units, resulting in parameters $\boldsymbol{\theta}\in\mathbb{R}^{1248}$. Our indirect methods use $S=4$ shooting segments. All trajectory optimization methods employ moving horizon optimization (horizon $H=3$) and 15 iterations per timestep.
@@
The dynamics model architecture is consistent with the previous Van der Pol setup, using ensemble neural ODEs ($\boldsymbol{\theta}\in\mathbb{R}^{1376}$) with identical control interpolation and ODE solver. Here, we use a shorter MPC horizon of $H=1$ with 25 iterations per time step, and two shooting segments ($S=2$) to ensure computational efficiency.
@@
We adopt an ensemble of neural ODEs with $M=5$ as a probabilistic dynamics model, consistent with the previous van der Pol setup ($\boldsymbol{\theta}\in\mathbb{R}^{1248}$) with identical control interpolation and ODE solver. We employ MPC with horizon $H=2$ and adjusted integration steps suitable for the Duffing oscillator’s faster time scale, and use $S=5$ shooting segments for the indirect methods.